# Consumption Partial Insurance

### v4

@@ 4.2 Estimation
We present estimates of the transmission parameters separately for the linear and quadratic specifications of the consumption function. For the linear function, we first present results when only second moments of income and consumption growth are targeted, exactly as in BPP. This purports to replicate BPP over the more recent sample period. We next report results for the linear function when, in addition to the second moments, third and fourth moments are also targeted. This helps assess whether and how, through the lenses of the linear specification, information in the tails of the income distribution affects the measurement of partial insurance. We then move on to the results from the quadratic specification. Our choice to estimate the income process in an earlier step enables the income parameters to remain unchanged regardless of the specification of the consumption function fitted. Appendix B.4 lists the moments targeted in each case.
Unless noted otherwise, we conduct inference for the main parameters using block bootstrap. We thus account for serial correlation and heteroskedasticity of arbitrary form and for parameter uncertainty from earlier steps of the estimation process. GMM (asymptotic) standard errors clustered at the household level are very similar, suggesting that earlier-step parameter uncertainty is negligible (the income parameters are tightly estimated). [^22] We compute bootstrap standard errors as the normalized interquartile range of bootstrap replications, i.e., the difference between the $75^{\text{th}}$ and $25^{\text{th}}$ percentiles divided by the interquartile range of the standard normal cdf. This is equivalent to applying a normal approximation to the distribution of bootstrap replications, thereby shielding the standard errors from extreme resampling draws (Blundell et al., 2016; Theloudis, 2021). Bootstrap standard errors without this approximation are similar. For the subsample analysis (tables 5-6 and corresponding tables in the appendix), we calculate GMM (asymptotic) standard errors clustered at the household level to avoid numerical instability in bootstrap, driven by vanishing shares of certain subsamples, e.g., the wealthy non-college graduates, in many bootstrap draws. For the replication of BPP using the original BPP data (table C.5), we show non-clustered GMM standard errors, as in BPP.
@@ 4.3 Empirical moments of income and consumption
Guvenen et al. (2021) focus on male earnings while our model and statistics for income growth above revolve around household disposable income. To facilitate the comparison, appendix C.2 reports moments of male earnings growth in our sample. The variance is similar between the PSID and the administrative data, at $0.271$ and $0.323$ respectively. Skewness is $-0.569$ (and highly statistically significant) versus $-1.039$ in the administrative data. While the nature of skewness is similar in both cases, the population data feature substantially more left skewness than the PSID; Guvenen et al. (2021) confirm this. Kurtosis, finally, is nearly identical between the two, at $13.711$ and $13.494$ respectively.
Left skewness and excess kurtosis are thus not just features of the growth rates in disposable household income (table 1), but they are also observed in alternative measures of income, such as male earnings (table C.2, panel A) and household earnings (table C.2, panel B). Together, left skewness and excess kurtosis suggest that income changes for most households are either close to their mean value of zero or far out in the left tail (large income reductions). These results corroborate (and, in fact, are close numerically with) findings by De Nardi et al. (2019); De Nardi et al. (2021) over an earlier period of the PSID.
@@
The moments of imputed consumption growth are in column 2 of table 1. The variance at $0.323$ is two-and-a-half times higher than the variance in the true series (column 1). In fact, the variance of imputed consumption growth is more than double the variance of income growth. The higher-order moments differ markedly between the imputed and true series. Skewness in imputed consumption growth is $-0.177$ (versus $-0.039$ in the true series), while kurtosis is $17.109$, almost four times higher than kurtosis in the true series. The imputation thus seems to impart substantial measurement/imputation error to consumption.
@@ 5.1 Income process
Targeting the third moments of income growth (in addition to the second moments) leaves the variances unchanged. We estimate the skewness (third standardized moment) at $-0.947$ and $-1.31$ for permanent and transitory shocks, respectively. Both shocks thus exhibit substantial negative skewness, i.e., their distribution has a longer left than right tail. This is consistent with Geweke and Keane (2000) who find left skewness of earnings shocks in the early PSID, and Theloudis (2021) who finds left skewness of wage shocks in the recent period. These results vary with the level of income, but we postpone this discussion for later. Assuming for now that good and bad shocks transmit similarly into consumption, the long left tail implies that bad shocks are more damaging than good ones because they are further away from the zero mean – they are more extreme. This is true for both permanent and transitory shocks. Skewness of income growth, however, is driven mostly by permanent shocks (fully so under stationarity). Therefore, if targeting the third moments of income growth subsequently changes the degree of consumption insurance in the household, this must reflect the implications of permanent rather than of transitory shocks.
@@ 5.2 Linear consumption function
What explains the discrepancy in $\phi^{(1)}$ with BPP? As suggested earlier, there are a few important differences between our replication and the original exercise in BPP: new consumption data and a different time period. To help decipher their role, we impute consumption from the CEX into the modern PSID in the same way BPP did for the earlier period, and we repeat the estimation. The results are in column 2 of table 3. The pass-through of permanent shocks doubles to $\phi^{(1)}=0.288$, while the pass-through of transitory shocks remains statistically zero (but at a larger magnitude). Using imputed consumption data thus inflates the transmission of permanent shocks and lowers the degree of insurance. This is likely the byproduct of measurement/imputation error, evident in the variance $\sigma^{2}_{u_{c}}$ of consumption error that more than triples between the true and the imputed data. By contrast, the variance of taste heterogeneity $\sigma^{2}_{\xi}$ remains similar in both cases, and very close to BPP. The imputation, therefore, seems to bring in large amounts of measurement error, explaining about 1/3 of the difference between our lower $\phi^{(1)}$ and BPP’s higher estimate. Appendix C.4 discusses the details of the imputation and a mechanism through which it inflates $\phi^{(1)}$ in particular (the imputation involves the inversion of a demand system, so certain types of measurement error may inflate the imputed income-consumption covariance – evident also in table 1).
@@
Column 3 of table 3 presents the results from targeting third- in addition to second-order moments of consumption and income. The transmission parameters remain largely unchanged upon inclusion of the third moments. We estimate $\phi^{(1)}=0.160$ (as opposed to $0.152$ with only second moments) and $\psi^{(1)}=-0.012$ (again indistinguishable from zero). We thus again find large partial insurance to permanent shocks and full insurance to transitory shocks. Skewness in taste heterogeneity and consumption error is small (much smaller than skewness in income shocks), while their variances remain unchanged from the specification with only second moments. Therefore, left skewness in consumption is driven by negatively skewed income shocks rather than by taste heterogeneity or measurement error.
@@
Column 4 of table 3 presents the results from targeting all income and consumption moments up to fourth-order. The earlier findings hold up. We estimate $\phi^{(1)}=0.112$ (versus $0.152$ and $0.160$ in the lower-order cases) and $\psi^{(1)}=0.007$ (indistinguishable from zero). A $10\%$ permanent change in income thus results in $1.12\%$ shift in consumption. Therefore, if anything, we find evidence for slightly more insurance to permanent shocks upon inclusion of the fourth moments. The dispersion and skewness of taste heterogeneity and consumption error do not change much from the lower-order cases; we find evidence for excess kurtosis in both cases, albeit at a much lesser extent than for income shocks. [^33] Appendix table C.8 uses imputed consumption from the CEX to target higher-order moments as in columns 3-4 of table 3. We do see slightly higher pass-through of permanent shocks upon inclusion of the higher-order moments, but we also find a strong, significant, and negative response to transitory shocks, which is hard to explain. We do not think those results are very credible; the imputation brings in substantial measurement/imputation error, which becomes worse the higher the order of the target moments is.
@@
Clearly, targeting higher-order moments in the context of the linear consumption function has not altered much the degree of partial insurance vis-à-vis targeting second-order moments alone. This is counterintuitive given the extent of tail income risk in the data, as well as the ample evidence in the literature pointing to the contrary. [^34] Ai and Bhandari (2021) show that tail risk is uninsured in realistically parameterized labor markets. However, De Nardi et al. (2019) and Busch and Ludwig (2024) find higher insurance when shocks are non-normal, due to stronger precautionary motives. The latter caution against BPP’s linear framework, showing evidence of bias in the presence of tail risk. Our quadratic specification addresses their concern. One would expect that tail risk increases the pass-through of shocks, as there must be a limit beyond which households cannot further insure consumption against extreme events. Yet, the linear consumption function leaves no room to such events. BPP rely on a log-linearization of the consumption rule around a deterministic lifecycle profile (appendix A). But tail events move the agent far from that profile: if utility is sufficiently concave, a large shock will command a large consumption response. A linearized marginal utility fails to pick up such large response (Carroll, 2001), which in turn explains why we do not see much action upon including higher-order moments. In other words, tail risk has, by design, no role in the linear consumption function, which in turn motivates our higher-order specification subsequently.
@@ 5.3.1 Main results
We estimate the transmission parameter of the average permanent shock (the coefficient on the linear part of a polynomial in the permanent shock) at $\phi^{(1)}=0.134$, close to the values 0.110.110.110.110.160.160.160.16 obtained from the linear specification. We estimate the transmission parameter of the average transitory shock (the coefficient on the linear part of a polynomial in the transitory shock) at $\psi^{(1)}=-0.003$, statistically indistinguishable from zero. As $\phi^{(1)}$ and $\psi^{(1)}$ remain largely unaltered relative to the linear case, interest lies in the transmission parameters of the typical shocks away from the middle of the distribution, i.e., the coefficients on the quadratic part of the polynomial. We estimate a negative $\phi^{(2)}=-0.04$ for permanent shocks and a positive (but statistically insignificant for now) $\psi^{(2)}=0.015$ for transitory shocks. The coefficient $\omega^{(22)}$ on the product of the two shocks is large but very imprecisely estimated. The variance of taste heterogeneity $\sigma_{\xi}^{2}$ and consumption measurement error $\sigma_{u_{c}}^{2}$ remain effectively unchanged from the linear case. [^37] Skewness and kurtosis of taste heterogeneity and consumption error are not identified because they require higher-order moments of consumption, which depend on higher than fourth-order moments of income.
The values of $\phi^{(2)}$ and $\psi^{(2)}$ have important implications for the pass-through of shocks into consumption. To help interpret the results, recall that the pass-through of permanent shocks is given by (6) and depends on $\phi^{(1)}$, $\phi^{(2)}$, $\omega^{(22)}$, and the magnitudes and signs of both income shocks. If, for simplicity, we set the transitory shock to its mean value ($v_{it}=0$) and plug in the parameter estimates from table 3, we get
@@
Therefore, bad permanent shocks are not just more extreme than good shocks, i.e., further away from the zero mean – see left skewness in table 2, but they also have larger pass-through rates into consumption. Bad permanent shocks are thus double hurtful.
Second, severe bad shocks are more hurtful than moderate ones because they too are associated with larger transmission parameters into consumption. Contrast, for example, a very large bad permanent shock, $\zeta_{it}=-0.7$, about four standard deviations away from the mean, with a moderate one, $\zeta_{it}=-0.07$, half standard deviation from the mean. In the first case, $\partial\Delta c_{it}/\partial\zeta_{it}=0.19$ (s.e.$=0.044$) while in the latter case $\partial\Delta c_{it}/\partial\zeta_{it}=0.139$ (s.e.$=0.026$). We can generalize this to bad shocks of any magnitude and conclude that
@@
Why would a positive or small shock be better insured? When a good shock hits, households save part of it and consumption does not respond too much. When a bad shock hits, however, households insure themselves by running down savings or by using other formal or informal insurance arrangements. But there is clearly a limit to how much they can dissave or borrow. As the magnitude of the bad shock worsens, these insurance arrangements will be eventually exhausted, resulting in a larger fraction of the shock transmitting into consumption. A large negative shock thus affects consumption by relatively more than a small or positive shock does. [^38] The results also imply that the pass-through of a good shock decreases with its magnitude. The reason behind this counterintuitive result is that we lack sufficient degrees of freedom to separate among all combinations of large/small, positive/negative shocks. However, large left skewness and excess kurtosis imply that extreme good shocks are rare. We should thus not extrapolate the implications of the parameter estimates for the case of extreme good shocks, as such shocks are effectively not present in our sample.
The asymmetric response to bad versus good income news has recently been documented in survey questions in which subjects directly report how they would adjust consumption in different hypothetical scenarios. For example, Bunn et al. (2018) find that the marginal propensity to consume from negative income shocks is larger than from positive shocks, while Christelis et al. (2019) and Fuster et al. (2020) find that responses to losses are much larger than responses to gains, and the magnitude of the response increases with the magnitude of the bad shock. This literature, however, elicits responses to transitory shocks. By contrast, we establish these asymmetries also in response to permanent shocks in a context in which the shocks are not directly observed (namely in survey data).
@@
Unlike the transmission parameters of permanent shocks, neither $\psi^{(1)}$ nor $\psi^{(2)}$ are statistically significant in the full sample. This implies that transitory shocks, be they good or bad, moderate or extreme, do not pass through into consumption. This is similar to our findings for the linear case as well as to BPP. By contrast, this is not the case for certain well-defined subsamples, which we analyze in the next subsections.
Despite the lack of statistical significance, the transmission parameters of transitory shocks have, at face value, an interesting implication: good and large shocks have a numerically larger pass-through than smaller ones, whose pass-through is economically zero. While we should be cautious in drawing conclusions about large good transitory shocks (given left skewness and excess kurtosis, such shocks are rare in our sample), this implication helps reconcile the large pass-through found in quasi-experimental studies with the negligible pass-through found in BPP and similar studies that employ covariance restrictions on survey data. The former typically exploit large infrequent transitory episodes (e.g., tax rebates or lotteries as in Parker et al., 2013; Misra and Surico, 2014; Fagereng et al., 2021), which are unlikely to be present in the pooled survey data that the latter studies exploit (e.g., BPP, Blundell et al., 2016; Theloudis, 2021). In an environment where the transmission of shocks depends on their magnitude, these two types of settings will lead to opposite conclusions. Our nonlinear consumption function reconciles the two by permitting transitory shocks to pass through to consumption at a varying rate that depends on the shocks’ magnitude and sign.
To summarize our findings from the quadratic consumption function: 1.) bad permanent shocks reduce consumption by more than equal-sized good permanent shocks increase it; 2.) severe bad permanent shocks reduce consumption by more than moderate ones do. For a given value of $\phi^{(1)}$, the econometrician will overestimate the true degree of partial insurance against bad permanent shocks if she assumes a linear consumption function instead of the nonlinear specification herein. By contrast, transitory shocks, be they good or bad, small or large, do not affect consumption, at least not in a statistically significant way. [^39] Similar conclusions are obtained if we use imputed consumption from the CEX. In particular, $\phi^{(2)}<0$, with a magnitude similar to that obtained from the PSID, and $\psi^{(2)}\approx>0$ (appendix table C.8).

### v5

@@ 4.2 Estimation
We present estimates of the transmission parameters separately for the linear and quadratic specifications of the consumption function. For the linear function, we first present results when only second moments of income and consumption growth are targeted, exactly as in BPP. This purports to replicate BPP over the more recent sample period. We next report results for the linear function when, in addition to the second moments, third and fourth moments are also targeted. This helps assess whether and how, through the lenses of the linear specification, information in the tails of the income distribution affects the measurement of partial insurance. We then move on to the results from the quadratic specification. Our choice to estimate the income process separately in an earlier step enables the income parameters to remain unchanged regardless of the specification of the consumption function fitted. Appendix B.4 lists the moments targeted in each case.
Unless noted otherwise, we conduct inference for the main parameters using block bootstrap. We thus account for serial correlation and heteroskedasticity of arbitrary form and for parameter uncertainty from earlier steps of the estimation process. GMM (asymptotic) standard errors clustered at the household level are very similar, suggesting that earlier-step parameter uncertainty is negligible (the income parameters are tightly estimated). [^23] We compute bootstrap standard errors as the normalized interquartile range of bootstrap replications, i.e., the difference between the $75^{\text{th}}$ and $25^{\text{th}}$ percentiles divided by the interquartile range of the standard normal cdf. This is equivalent to applying a normal approximation to the distribution of bootstrap replications, thereby shielding the standard errors from extreme resampling draws (Blundell et al., 2016; Theloudis, 2021). Bootstrap standard errors without this approximation are similar. For the subsample analysis (tables 5-6 and corresponding tables in the appendix), we calculate GMM (asymptotic) standard errors clustered at the household level to avoid numerical instability in bootstrap, driven by vanishing shares of certain subsamples, e.g., the wealthy non-college graduates, in many bootstrap draws. For the replication of BPP using the original BPP data (table C.5), we show non-clustered GMM standard errors as in BPP.
@@ 4.3 Empirical moments of income and consumption
Guvenen et al. (2021) focus on male earnings while our model and statistics for income growth above revolve around household disposable income. To facilitate a comparison, appendix C.2 reports moments of male earnings growth in our sample. The variance is similar between the PSID and the administrative data, at $0.271$ and $0.323$ respectively. Skewness is $-0.569$ (and highly statistically significant) versus $-1.039$ in the administrative data. While the nature of skewness is similar in both cases, the population data feature substantially more left skewness than the PSID; Guvenen et al. (2021) also confirmed this. Kurtosis, finally, is nearly identical between the two, at $13.711$ and $13.494$ respectively.
Left skewness and excess kurtosis are thus not just features of the growth rates in disposable household income (table 1), but they are also observed in alternative measures of income, such as male earnings (table C.2, panel A) and household earnings (table C.2, panel B). Together, left skewness and excess kurtosis suggest that income changes for most households are either close to their mean value of zero or far out in the left tail (large income reductions). [^27] Using earnings as the relevant income variable in the model does not materially change our subsequent results, in particular the asymmetry in the transmission of bad versus good (earnings) shocks. These results corroborate (and, in fact, are close numerically with) findings by De Nardi et al. (2019); De Nardi et al. (2021) over an earlier period of the PSID.
@@
The moments of imputed consumption growth are in column 2 of table 1. The variance at $0.323$ is two-and-a-half times larger than the variance in the true series (column 1). In fact, the variance of imputed consumption growth is more than double the variance of income growth. The magnitude of the higher-order moments is also multiple times larger in the imputed versus the true series. Skewness in imputed consumption growth is $-0.177$ (versus $-0.039$ in the true series), while kurtosis is $17.109$, almost four times higher than kurtosis in the true series. The imputation thus seems to impart substantial measurement/imputation error to consumption.
@@ 5.1 Income process
Targeting the third moments of income growth (in addition to the second moments) leaves the variances unchanged. We estimate the skewness (third standardized moment) at $-0.947$ and $-1.31$ for permanent and transitory shocks, respectively. Both shocks thus exhibit substantial negative skewness, i.e., their distribution has a longer left than right tail. This is consistent with Geweke and Keane (2000) who find left skewness of earnings shocks in the early PSID, and Theloudis (2021) who finds left skewness of wage shocks in the recent period. These results vary with the level of income, but we postpone that discussion for later. Assuming for now that good and bad shocks transmit similarly into consumption, the long left tail implies that bad shocks are more damaging than good ones because they are farther away from the zero mean – they are more extreme. This is true for both permanent and transitory shocks. Skewness of income growth, however, is driven mostly by permanent shocks (fully so under stationarity). Therefore, if targeting the third moments of income growth subsequently changes the degree of consumption insurance in the household, this must reflect the implications of permanent rather than of transitory shocks.
@@ 5.2 Linear consumption function
What explains the discrepancy in $\phi^{(1)}$ with BPP? As suggested earlier, there are a few important differences between our replication and the original exercise in BPP: new consumption data and a different time period. To help decipher their role, we impute consumption from the CEX into the modern PSID in the same way BPP did for the earlier period, and we repeat the estimation. The results are in column 2 of table 3. The pass-through of permanent shocks doubles to $\phi^{(1)}=0.288$, while the pass-through of transitory shocks remains statistically zero (but at a larger magnitude). Using imputed consumption data thus inflates the transmission of permanent shocks and lowers the degree of insurance. This is likely the byproduct of measurement/imputation error, evident in the variance $\sigma^{2}_{u_{c}}$ of consumption error that more than triples between the true and the imputed data. By contrast, the variance of taste heterogeneity $\sigma^{2}_{\xi}$ remains similar in both cases, and very close to BPP. The imputation, therefore, seems to bring in large amounts of measurement error, explaining about 1/3 of the difference between our lower $\phi^{(1)}$ and BPP’s higher estimate. Appendix C.4 discusses the details of the imputation and a mechanism through which in particular it inflates $\phi^{(1)}$: the imputation involves the inversion of a demand system, so certain types of measurement error can inflate the imputed income-consumption covariance – evident also in table 1.
@@
Column 3 of table 3 presents the results from targeting third- in addition to second-order moments of consumption and income in the PSID. The transmission parameters remain largely unchanged. We estimate $\phi^{(1)}=0.160$ (as opposed to $0.152$ with only second moments) and $\psi^{(1)}=-0.012$ (again indistinguishable from zero). We thus again find large partial insurance to permanent shocks and full insurance to transitory shocks. Skewness in taste heterogeneity and consumption error is small (much smaller than skewness in income shocks), while their variances remain unchanged from the specification with only second moments. Therefore, left skewness in consumption is driven by negatively skewed income shocks rather than by taste heterogeneity or measurement error.
@@
Column 4 of table 3 presents the results from targeting all income and consumption moments up to fourth-order. The earlier findings hold up. We estimate $\phi^{(1)}=0.112$ (versus $0.152$ and $0.160$ in the lower-order cases) and $\psi^{(1)}=0.007$ (indistinguishable from zero). A $10\%$ permanent change in income thus results in $1.12\%$ shift in consumption. Therefore, if anything, we find evidence for slightly more insurance to permanent shocks upon inclusion of the fourth moments. The dispersion and skewness of taste heterogeneity and consumption error do not change much from the lower-order cases; we find evidence for excess kurtosis in both cases, albeit at a much lesser extent than for income shocks. [^35] Appendix table C.8 uses imputed consumption from the CEX to target higher-order moments as in columns 3-4 of table 3. We do see slightly higher pass-through of permanent shocks upon inclusion of the higher-order moments, but we also find a strong negative response to transitory shocks, which is hard to explain. We do not think those results are very credible; the imputation brings in substantial measurement and imputation error, which likely becomes worse the higher the order of the target moments is.
@@
Clearly, targeting higher-order moments within the linear consumption function has not altered much the degree of partial insurance vis-à-vis targeting second-order moments alone. This is counterintuitive given the extent of tail income risk in the data and ample evidence in the literature to the contrary. [^36] Ai and Bhandari (2021) show that tail risk is uninsured in realistically parameterized labor markets. However, De Nardi et al. (2019) and Busch and Ludwig (2024) find higher insurance when shocks are non-normal, due to stronger precautionary motives. The latter caution against BPP’s linear framework, showing evidence of bias in the presence of tail risk. Our quadratic specification addresses their concern. One would expect that tail risk increases the pass-through of shocks, as there must be a limit beyond which households cannot further insure consumption against extreme events, even if such risk triggers a stronger precautionary motive. Yet, the linear consumption function leaves no room to such events, because it reflects a linearization of the consumption rule around a deterministic lifecycle profile. Tail events, however, move the agent far from that profile: if utility is sufficiently concave, a large shock will command a large consumption response. A linearized marginal utility fails to pick up such large response, which in turn explains why we do not see much action upon including higher-order moments. In other words, tail risk has, by design, no role in the linear consumption function, which in turn motivates our higher-order specification subsequently.
@@ 5.3.1 Main results
We estimate the transmission parameter of the average permanent shock (the coefficient on the linear part of the consumption polynomial) at $\phi^{(1)}=0.134$, close to the values 0.110.110.110.110.160.160.160.16 obtained from the linear specification. We estimate the transmission parameter of the average transitory shock at $\psi^{(1)}=-0.003$, statistically indistinguishable from zero. As $\phi^{(1)}$ and $\psi^{(1)}$ remain largely unaltered relative to the linear case, interest lies in the transmission parameters of the typical shocks away from the middle of the distribution, i.e., the coefficients on the quadratic part of the polynomial.
We estimate a negative $\phi^{(2)}=-0.04$ for permanent shocks but a positive $\psi^{(2)}=0.015$ for transitory ones. $\phi^{(2)}$ is statistically insignificant at the margin for now but turns highly significant subsequently. By contrast, $\psi^{(2)}$ is almost always insignificant, barring the few exceptions we highlight below. The coefficient $\omega^{(22)}$ on the product of the two shocks is large but always imprecise. The variance of taste heterogeneity $\sigma_{\xi}^{2}$ and consumption measurement error $\sigma_{u_{c}}^{2}$ remain effectively unchanged from the linear case. [^39] Skewness and kurtosis of taste heterogeneity and consumption error are not identified because they require higher-order moments of consumption, which depend on higher than fourth-order moments of income.
At face value, $\phi^{(2)}$ has important implications for the pass-through of permanent shocks into consumption. To help interpret the results, recall that the pass-through is given by (6) and depends on $\phi^{(1)}$, $\phi^{(2)}$, $\omega^{(22)}$, and the magnitudes and signs of both income shocks. If, for simplicity, we set the transitory shock to its mean value ($v_{it}=0$) and plug in the parameter estimates from table 3, we get
@@
Therefore, bad permanent shocks are not just more extreme than good shocks, i.e., farther away from the zero mean – see left skewness in table 2, but they also have larger pass-through rates into consumption. Bad permanent shocks are thus double hurtful.
Second, severe bad shocks are more hurtful than moderate ones because they too are associated with larger transmission parameters into consumption. Compare, for example, a very large bad permanent shock, $\zeta_{it}=-0.7$, about four standard deviations away from the mean, with a moderate one, $\zeta_{it}=-0.07$, half standard deviation from the mean. In the first case, $\partial\Delta c_{it}/\partial\zeta_{it}=0.19$ (s.e.$=0.044$) while in the latter case $\partial\Delta c_{it}/\partial\zeta_{it}=0.139$ (s.e.$=0.026$). We can generalize this to bad shocks of any magnitude and conclude that
@@
Why would a positive or small shock be better insured? When a good shock hits, households save part of it and consumption does not respond too much. When a bad shock hits, however, households insure themselves by running down savings or by using other formal or informal insurance arrangements. Yet there is clearly a limit to how much they can dissave or borrow. As the magnitude of the bad shock worsens, these insurance arrangements will be eventually exhausted, resulting in a larger fraction of the shock transmitting into consumption. A large negative shock thus affects consumption by relatively more than a small or positive shock does. [^40] The results also imply that the pass-through of a good shock decreases with its magnitude. The reason behind this counterintuitive result is that we lack sufficient degrees of freedom to separate among all combinations of large/small, positive/negative shocks. However, large left skewness and excess kurtosis imply that extreme good shocks are rare. We should thus not extrapolate the implications of the parameter estimates for the case of extreme good shocks, as such shocks are effectively not present in our sample.
The asymmetric response to bad versus good income news has recently been documented in survey questions in which subjects directly report how they would adjust consumption in different hypothetical scenarios. For example, Bunn et al. (2018) find that the marginal propensity to consume from negative income shocks is larger than from positive shocks, while Christelis et al. (2019) and Fuster et al. (2020) find that responses to losses are much larger than responses to gains, and the magnitude of the response increases with the magnitude of the bad shock. This literature, however, elicits responses to transitory shocks. By contrast, we establish these asymmetries in response to permanent shocks in a context in which the shocks are not directly observed (namely in survey data).
@@
Unlike the transmission parameters of permanent shocks, neither $\psi^{(1)}$ nor $\psi^{(2)}$ are statistically significant in the full sample. This implies that transitory shocks, be they good or bad, moderate or extreme, do not pass through into consumption.
Despite their statistical insignificance, the transmission parameters of transitory shocks have, at face value, an interesting implication: good and large shocks have a numerically larger pass-through than smaller ones, whose pass-through is economically zero. While we should be cautious in drawing conclusions about large good transitory shocks (given left skewness and excess kurtosis, such shocks are rare in our sample), this implication helps reconcile the large pass-through found in quasi-experimental studies with the negligible pass-through found in BPP and similar studies that employ covariance restrictions on survey data. The former typically exploit large infrequent transitory gains (e.g., tax rebates or lotteries as in Parker et al., 2013; Misra and Surico, 2014; Fagereng et al., 2021), which are unlikely to be present in the pooled survey data that the latter studies exploit (e.g., BPP, Blundell et al., 2016; Theloudis, 2021). In an environment where the transmission of shocks depends on their magnitude, these two types of settings will lead to opposite conclusions. Our framework can potentially reconcile the two by permitting transitory shocks to pass through to consumption at a varying rate that depends on the shocks’ magnitude and sign.
To summarize our findings from the quadratic function up to this point: 1.) bad permanent shocks reduce consumption by more than equal-sized good permanent shocks increase it; 2.) severe bad permanent shocks reduce consumption by more than moderate ones do. For a given value of $\phi^{(1)}$, the econometrician will overestimate the true degree of partial insurance against bad permanent shocks if she assumes a linear consumption function instead of the nonlinear specification herein. By contrast, transitory shocks, be they good or bad, small or large, do not affect consumption, at least not in a statistically significant way. [^41] Similar conclusions are obtained if we use imputed consumption from the CEX. In particular, $\phi^{(2)}<0$, with a magnitude similar to that obtained from the PSID, and $\psi^{(2)}\approx>0$ (appendix table C.8).